Two pumps of different sizes, working together, can empty a tank in 7 hours. The larger pump can empty the tank in 4 hours less than the smaller pump. If the larger pump is out of order, how long will it take the smaller pump to complete the job?

Answer by richwmiller(17219)

7/x+7/(x-4)=1
1/53 (x-9)^2 = 1
x = 1.7199
x = 16.280
It looks like there are two positive answers but remember that we must subtract 4 to get the smaller pumps time.
x = 16.280
x-4 = 12.280